It is found that, in general 2 out of 5 persons are swimmer.
If 4 persons are selected at random, what is the probability that -
(i) two of them are swimmers and
(ii) at least on of them is a swimmer?

Answer by Finavon(81)

Each selected from general population has p(swim) 0.4; p(non) 0.6
a)
SSNNN
SNSNN
SNNSN
SNNNS
NSSNN
NSNSN
NSNNS
NNSSN
NNSNS
NNNSS
P Each 0.6^3*0.4^2=0.03456
10 ways so x10
p(2swim)=0.3456
b) need Prob not all non-swimmers
= (1-(0.6*0.6*0.6*0.6)) = 0.87